Find the sum of the indicated terms of the geometric sequence: 6, 18, 54 . . . (to 7 terms)

Found 2 solutions by JBarnum, camerontdy:
Answer by JBarnum(2146)

6*3=18
18*3=54
54*3=162
162*3=486
486*3=1458
1458*3=4374
6+18+54+162+486+1458+4374=6540

Answer by camerontdy(1)

6*3=18
18*3=54
54*3=162
162*3=486
486*3=1458
1458*3=4374
6+18+54+162+486+1458+4374=6558